# Deep Modularity Networks with Diversity–Preserving Regularization

### v1

@@ B.3 PubMed
993 DMoN 10.64 55.82 21.92 39.57
550 DMoN 7.91 58.30 26.08 48.88
243 DMoN 7.87 57.66 21.90 44.67
16 DMoN 8.02 57.40 21.90 44.83
716 DMoN 8.02 57.35 21.98 45.16
383 DMoN 12.56 53.65 26.03 49.14
277 DMoN 7.78 57.61 21.73 44.90
274 DMoN 7.94 57.44 21.59 44.78
188 DMoN 7.90 57.50 21.86 45.03
796 DMoN 7.93 57.53 21.76 44.82
Mean 8.66 57.03 22.68 45.18
Variance 0.03 0.02 0.03 0.07
@@
993 DMoN–DPR (D) 10.63 55.83 21.92 39.59
550 DMoN–DPR (D) 7.94 58.27 26.16 48.94
243 DMoN–DPR (D) 7.88 57.65 21.93 44.67
16 DMoN–DPR (D) 8.00 57.41 21.80 44.75
716 DMoN–DPR (D) 8.07 57.30 21.97 45.11
383 DMoN–DPR (D) 12.53 53.69 25.84 49.03
277 DMoN–DPR (D) 7.79 57.60 21.73 44.91
274 DMoN–DPR (D) 7.96 57.43 21.76 44.94
188 DMoN–DPR (D) 7.91 57.49 21.87 45.05
796 DMoN–DPR (D) 7.92 57.54 21.84 44.89
Mean 8.66 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
@@
993 DMoN–DPR (V) 10.64 55.82 21.91 39.57
550 DMoN–DPR (V) 7.92 58.29 26.08 48.88
243 DMoN–DPR (V) 7.88 57.65 21.91 44.68
16 DMoN–DPR (V) 8.01 57.40 21.92 44.86
716 DMoN–DPR (V) 8.02 57.35 21.97 45.16
383 DMoN–DPR (V) 12.56 53.64 26.03 49.14
277 DMoN–DPR (V) 7.78 57.61 21.75 44.89
274 DMoN–DPR (V) 7.94 57.44 21.59 44.78
188 DMoN–DPR (V) 7.91 57.49 21.86 45.04
796 DMoN–DPR (V) 7.92 57.54 21.78 44.83
Mean 8.66 57.02 22.68 45.18
Variance 0.03 0.02 0.03 0.07
@@
993 DMoN–DPR (E) 10.64 55.82 21.91 39.59
550 DMoN–DPR (E) 7.92 58.29 26.08 48.88
243 DMoN–DPR (E) 7.88 57.65 21.90 44.67
16 DMoN–DPR (E) 8.01 57.40 21.90 44.83
716 DMoN–DPR (E) 8.03 57.34 21.96 45.13
383 DMoN–DPR (E) 12.55 53.65 26.04 49.15
277 DMoN–DPR (E) 7.78 57.61 21.75 44.91
274 DMoN–DPR (E) 7.96 57.42 21.63 44.80
188 DMoN–DPR (E) 7.93 57.47 21.88 45.06
796 DMoN–DPR (E) 7.95 57.52 21.79 44.85
Mean 8.67 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
@@
993 DMoN–DPR (DV) 10.63 55.84 21.89 39.58
550 DMoN–DPR (DV) 7.93 58.27 26.15 48.92
243 DMoN–DPR (DV) 7.89 57.64 21.94 44.67
16 DMoN–DPR (DV) 7.97 57.43 21.83 44.78
716 DMoN–DPR (DV) 8.06 57.31 22.00 45.11
383 DMoN–DPR (DV) 12.52 53.69 25.81 49.01
277 DMoN–DPR (DV) 7.79 57.59 21.72 44.91
274 DMoN–DPR (DV) 7.93 57.46 21.73 44.92
188 DMoN–DPR (DV) 7.91 57.48 21.83 45.03
796 DMoN–DPR (DV) 7.93 57.53 21.86 44.93
Mean 8.66 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
@@
993 DMoN–DPR (DE) 10.60 55.87 21.90 39.66
550 DMoN–DPR (DE) 7.94 58.27 26.15 48.91
243 DMoN–DPR (DE) 7.89 57.64 21.93 44.67
16 DMoN–DPR (DE) 7.97 57.43 21.83 44.78
716 DMoN–DPR (DE) 8.08 57.29 22.01 45.14
383 DMoN–DPR (DE) 12.54 53.68 25.80 49.00
277 DMoN–DPR (DE) 7.79 57.59 21.71 44.90
274 DMoN–DPR (DE) 7.97 57.42 21.75 44.94
188 DMoN–DPR (DE) 7.93 57.47 21.86 45.01
796 DMoN–DPR (DE) 7.92 57.54 21.78 44.87
Mean 8.66 57.02 22.67 45.19
Variance 0.03 0.02 0.03 0.07
@@
993 DMoN–DPR (VE) 10.64 55.83 21.89 39.57
550 DMoN–DPR (VE) 7.92 58.29 26.09 48.88
243 DMoN–DPR (VE) 7.87 57.65 21.89 44.67
16 DMoN–DPR (VE) 8.01 57.40 21.91 44.85
716 DMoN–DPR (VE) 8.03 57.33 21.96 45.13
383 DMoN–DPR (VE) 12.57 53.63 26.02 49.14
277 DMoN–DPR (VE) 7.78 57.60 21.74 44.91
274 DMoN–DPR (VE) 7.96 57.42 21.61 44.81
188 DMoN–DPR (VE) 7.92 57.47 21.90 45.03
796 DMoN–DPR (VE) 7.94 57.52 21.77 44.84
Mean 8.66 57.01 22.68 45.18
Variance 0.03 0.02 0.03 0.07
@@
993 DMoN–DPR (DVE) 10.59 55.87 21.90 39.65
550 DMoN–DPR (DVE) 7.93 58.27 26.14 48.91
243 DMoN–DPR (DVE) 7.89 57.64 21.93 44.67
16 DMoN–DPR (DVE) 7.97 57.42 21.84 44.78
716 DMoN–DPR (DVE) 8.08 57.29 21.91 45.06
383 DMoN–DPR (DVE) 12.51 53.69 25.79 48.98
277 DMoN–DPR (DVE) 7.79 57.58 21.70 44.88
274 DMoN–DPR (DVE) 7.94 57.45 21.74 44.94
188 DMoN–DPR (DVE) 7.94 57.45 21.89 45.04
796 DMoN–DPR (DVE) 7.92 57.54 21.81 44.89
Mean 8.66 57.02 22.67 45.18
Variance 0.03 0.02 0.03 0.07
B.4 Coauthor CS
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor CS dataset.
@@ B.4 Coauthor CS
993 DMoN 19.07 72.23 68.48 58.27
550 DMoN 19.14 71.64 66.78 55.19
243 DMoN 18.11 72.88 70.78 60.88
16 DMoN 18.28 73.04 71.13 62.23
716 DMoN 18.41 72.63 69.20 59.44
383 DMoN 18.45 72.43 69.96 59.60
277 DMoN 19.53 72.25 67.51 54.79
274 DMoN 20.20 71.69 69.38 58.09
188 DMoN 17.89 73.05 71.52 60.61
796 DMoN 18.89 72.50 69.77 59.18
Mean 18.80 72.43 69.45 58.83
Variance 0.01 0.00 0.02 0.06
@@
993 DMoN–DPR(D) 18.88 72.53 72.11 62.71
550 DMoN–DPR(D) 20.80 71.19 68.18 57.76
243 DMoN–DPR(D) 19.36 71.83 71.19 62.87
16 DMoN–DPR(D) 18.99 72.42 71.12 63.30
716 DMoN–DPR(D) 17.85 73.08 70.14 56.54
383 DMoN–DPR(D) 19.24 71.60 69.23 60.49
277 DMoN–DPR(D) 18.20 73.02 72.35 64.04
274 DMoN–DPR(D) 18.86 72.11 70.97 61.56
188 DMoN–DPR(D) 19.85 71.18 68.61 59.59
796 DMoN–DPR(D) 18.02 72.72 71.70 62.08
Mean 19.01 72.17 70.56 61.09
Variance 0.01 0.00 0.02 0.06
@@
993 DMoN–DPR (V) 18.87 72.35 68.79 58.77
550 DMoN–DPR (V) 19.03 71.68 67.07 56.15
243 DMoN–DPR (V) 18.27 72.74 71.13 61.15
16 DMoN–DPR (V) 18.28 72.94 71.38 62.48
716 DMoN–DPR (V) 19.77 71.86 68.11 57.05
383 DMoN–DPR (V) 18.53 72.22 70.15 60.21
277 DMoN–DPR (V) 19.78 72.14 67.97 55.30
274 DMoN–DPR (V) 19.57 72.49 70.45 59.36
188 DMoN–DPR (V) 18.24 72.89 71.02 59.87
796 DMoN–DPR (V) 18.52 72.76 71.57 60.67
Mean 18.89 72.41 69.76 59.10
Variance 0.00 0.00 0.03 0.05
@@
993 DMoN–DPR (E) 18.95 71.61 72.94 59.97
550 DMoN–DPR (E) 19.87 71.23 70.05 56.24
243 DMoN–DPR (E) 20.13 70.45 71.39 60.47
16 DMoN–DPR (E) 19.44 71.06 72.47 64.02
716 DMoN–DPR (E) 19.11 71.28 73.55 64.42
383 DMoN–DPR (E) 20.09 70.20 67.61 54.84
277 DMoN–DPR (E) 21.27 69.71 68.97 56.84
274 DMoN–DPR (E) 20.01 71.45 73.13 62.46
188 DMoN–DPR (E) 18.70 71.68 72.18 63.95
796 DMoN–DPR (E) 19.74 71.04 71.06 60.77
Mean 19.73 70.97 71.34 60.40
Variance 0.01 0.00 0.04 0.12
@@
993 DMoN–DPR (DV) 18.76 72.48 72.77 63.41
550 DMoN–DPR (DV) 19.71 71.95 70.36 61.52
243 DMoN–DPR (DV) 19.10 72.09 71.75 63.65
16 DMoN–DPR (DV) 18.57 72.77 73.03 63.98
716 DMoN–DPR (DV) 19.02 72.54 70.44 58.08
383 DMoN–DPR (DV) 18.57 72.33 72.29 63.93
277 DMoN–DPR (DV) 17.94 73.14 72.46 63.13
274 DMoN–DPR (DV) 19.41 71.73 70.41 60.45
188 DMoN–DPR (DV) 20.33 71.18 69.03 58.99
796 DMoN–DPR (DV) 18.06 72.38 70.79 60.99
Mean 18.95 72.26 71.33 61.81
Variance 0.01 0.00 0.02 0.05
@@
993 DMoN–DPR (DE) 19.82 70.82 71.87 64.34
550 DMoN–DPR (DE) 19.72 70.51 70.06 61.80
243 DMoN–DPR (DE) 19.74 70.31 70.33 61.84
16 DMoN–DPR (DE) 19.83 71.73 74.31 63.08
716 DMoN–DPR (DE) 18.97 71.30 71.50 62.63
383 DMoN–DPR (DE) 19.97 70.74 72.47 65.26
277 DMoN–DPR (DE) 19.50 71.15 72.02 60.54
274 DMoN–DPR (DE) 19.42 69.84 68.51 59.93
188 DMoN–DPR (DE) 21.62 69.25 67.61 56.77
796 DMoN–DPR (DE) 20.19 69.93 69.69 61.41
Mean 19.88 70.56 70.84 61.76
Variance 0.00 0.01 0.04 0.06
@@
993 DMoN–DPR (VE) 19.08 71.49 73.11 60.17
550 DMoN–DPR (VE) 19.95 71.16 70.08 56.20
243 DMoN–DPR (VE) 20.21 70.35 71.26 60.33
16 DMoN–DPR (VE) 19.62 70.89 72.13 63.50
716 DMoN–DPR (VE) 19.16 71.25 73.42 64.30
383 DMoN–DPR (VE) 20.23 70.06 67.62 54.87
277 DMoN–DPR (VE) 20.14 70.81 70.43 59.06
274 DMoN–DPR (VE) 20.04 71.40 73.23 62.31
188 DMoN–DPR (VE) 18.72 71.67 72.17 63.93
796 DMoN–DPR (VE) 19.80 70.97 71.09 60.92
Mean 19.70 71.01 71.45 60.56
Variance 0.00 0.00 0.03 0.10
@@
993 DMoN–DPR (DVE) 19.32 71.13 72.76 65.12
550 DMoN–DPR (DVE) 19.75 70.50 70.00 61.78
243 DMoN–DPR (DVE) 19.97 70.19 70.40 61.86
16 DMoN–DPR (DVE) 18.87 72.23 74.23 65.17
716 DMoN–DPR (DVE) 19.03 71.26 71.56 62.76
383 DMoN–DPR (DVE) 20.06 70.65 72.47 65.18
277 DMoN–DPR (DVE) 19.06 72.07 73.97 64.36
274 DMoN–DPR (DVE) 19.41 69.81 68.57 60.03
188 DMoN–DPR (DVE) 21.66 69.20 67.65 56.86
796 DMoN–DPR (DVE) 20.30 69.78 69.21 60.98
Mean 19.74 70.68 71.08 62.41
Variance 0.01 0.01 0.05 0.07
B.5 Coauthor Physics
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor Physics dataset.

### v2

@@ E.3 PubMed
993 DMoN 10.94 55.54 22.49 40.20
550 DMoN 7.64 58.82 26.92 49.53
243 DMoN 7.93 57.62 21.77 44.56
16 DMoN 7.96 57.63 21.74 44.37
716 DMoN 8.05 57.34 21.70 44.97
383 DMoN 9.79 56.12 22.79 37.01
277 DMoN 7.80 57.63 21.69 44.89
274 DMoN 7.90 57.45 21.60 44.80
188 DMoN 10.09 55.60 21.39 37.09
796 DMoN 7.96 57.52 21.76 44.85
Mean $\pm$ Std 8.61 $\pm$ 1.19 57.13 $\pm$ 1.04 22.38 $\pm$ 1.65 43.23 $\pm$ 3.93
@@
993 DMoN–DPR (D) 10.82 55.67 22.35 40.01
550 DMoN–DPR (D) 7.65 58.81 26.89 49.57
243 DMoN–DPR (D) 7.95 57.60 21.79 44.56
16 DMoN–DPR (D) 7.99 57.59 21.73 44.35
716 DMoN–DPR (D) 8.05 57.36 21.71 44.95
383 DMoN–DPR (D) 9.78 56.13 22.83 37.08
277 DMoN–DPR (D) 7.81 57.62 21.67 44.88
274 DMoN–DPR (D) 7.82 57.53 21.67 44.91
188 DMoN–DPR (D) 10.14 55.54 21.34 37.09
796 DMoN–DPR (D) 7.97 57.51 21.82 44.96
Mean $\pm$ Std 8.60 $\pm$ 1.17 57.14 $\pm$ 1.03 22.38 $\pm$ 1.64 43.24 $\pm$ 3.95
@@
993 DMoN–DPR (V) 10.94 55.54 22.54 40.25
550 DMoN–DPR (V) 7.64 58.82 26.91 49.52
243 DMoN–DPR (V) 7.94 57.61 21.76 44.54
16 DMoN–DPR (V) 7.96 57.63 21.73 44.37
716 DMoN–DPR (V) 8.05 57.34 21.70 44.97
383 DMoN–DPR (V) 9.78 56.13 22.85 37.03
277 DMoN–DPR (V) 7.80 57.63 21.69 44.89
274 DMoN–DPR (V) 7.91 57.45 21.60 44.79
188 DMoN–DPR (V) 10.09 55.60 21.38 37.07
796 DMoN–DPR (V) 7.98 57.50 21.76 44.86
Mean $\pm$ Std 8.61 $\pm$ 1.19 57.12 $\pm$ 1.04 22.39 $\pm$ 1.65 43.23 $\pm$ 3.93
@@
993 DMoN–DPR (E) 10.96 55.52 22.51 40.19
550 DMoN–DPR (E) 7.65 58.81 26.90 49.50
243 DMoN–DPR (E) 7.94 57.61 21.76 44.55
16 DMoN–DPR (E) 7.97 57.62 21.74 44.37
716 DMoN–DPR (E) 8.05 57.33 21.63 44.91
383 DMoN–DPR (E) 9.82 56.08 22.79 36.95
277 DMoN–DPR (E) 7.80 57.63 21.69 44.89
274 DMoN–DPR (E) 7.91 57.45 21.60 44.80
188 DMoN–DPR (E) 10.08 55.56 21.26 37.16
796 DMoN–DPR (E) 7.98 57.51 21.75 44.85
Mean $\pm$ Std 8.62 $\pm$ 1.19 57.11 $\pm$ 1.05 22.36 $\pm$ 1.66 43.22 $\pm$ 3.92
@@
993 DMoN–DPR (DV) 10.84 55.65 22.38 40.02
550 DMoN–DPR (DV) 7.65 58.81 26.87 49.55
243 DMoN–DPR (DV) 7.95 57.60 21.79 44.56
16 DMoN–DPR (DV) 7.99 57.59 21.74 44.37
716 DMoN–DPR (DV) 8.06 57.35 21.72 44.96
383 DMoN–DPR (DV) 9.76 56.15 22.84 37.08
277 DMoN–DPR (DV) 7.81 57.62 21.67 44.88
274 DMoN–DPR (DV) 7.82 57.53 21.67 44.91
188 DMoN–DPR (DV) 10.12 55.55 21.30 37.10
796 DMoN–DPR (DV) 7.97 57.50 21.79 44.92
Mean $\pm$ Std 8.60 $\pm$ 1.17 57.13 $\pm$ 1.03 22.38 $\pm$ 1.64 43.23 $\pm$ 3.95
@@
993 DMoN–DPR (DE) 10.83 55.65 22.42 40.03
550 DMoN–DPR (DE) 7.65 58.80 26.87 49.55
243 DMoN–DPR (DE) 7.95 57.60 21.79 44.57
16 DMoN–DPR (DE) 8.02 57.57 21.69 44.34
716 DMoN–DPR (DE) 8.03 57.37 21.77 45.03
383 DMoN–DPR (DE) 9.81 56.10 22.81 37.07
277 DMoN–DPR (DE) 7.81 57.62 21.67 44.87
274 DMoN–DPR (DE) 7.82 57.52 21.68 44.92
188 DMoN–DPR (DE) 10.10 55.50 21.32 37.24
796 DMoN–DPR (DE) 7.98 57.50 21.79 44.92
Mean $\pm$ Std 8.60 $\pm$ 1.17 57.12 $\pm$ 1.04 22.38 $\pm$ 1.63 43.25 $\pm$ 3.93
@@
993 DMoN–DPR (VE) 10.85 55.64 22.48 40.10
550 DMoN–DPR (VE) 7.65 58.82 26.90 49.50
243 DMoN–DPR (VE) 7.94 57.61 21.76 44.54
16 DMoN–DPR (VE) 7.97 57.62 21.74 44.37
716 DMoN–DPR (VE) 8.06 57.33 21.63 44.91
383 DMoN–DPR (VE) 9.80 56.09 22.79 36.98
277 DMoN–DPR (VE) 7.80 57.63 21.68 44.88
274 DMoN–DPR (VE) 7.92 57.44 21.60 44.79
188 DMoN–DPR (VE) 10.01 55.65 21.43 37.15
796 DMoN–DPR (VE) 7.97 57.50 21.73 44.84
Mean $\pm$ Std 8.60 $\pm$ 1.16 57.13 $\pm$ 1.02 22.37 $\pm$ 1.65 43.21 $\pm$ 3.93
@@
993 DMoN–DPR (DVE) 10.85 55.63 22.40 40.04
550 DMoN–DPR (DVE) 7.66 58.80 26.86 49.54
243 DMoN–DPR (DVE) 7.94 57.61 21.79 44.57
16 DMoN–DPR (DVE) 8.02 57.57 21.69 44.34
716 DMoN–DPR (DVE) 8.07 57.34 21.68 44.93
383 DMoN–DPR (DVE) 9.81 56.09 22.81 37.08
277 DMoN–DPR (DVE) 7.81 57.62 21.67 44.87
274 DMoN–DPR (DVE) 7.83 57.51 21.69 44.93
188 DMoN–DPR (DVE) 10.15 55.47 21.31 37.20
796 DMoN–DPR (DVE) 7.97 57.50 21.75 44.87
Mean $\pm$ Std 8.61 $\pm$ 1.18 57.11 $\pm$ 1.05 22.37 $\pm$ 1.64 43.24 $\pm$ 3.92
E.4 Coauthor CS
The following tables list the performance of DiffPool, MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the Coauthor CS dataset.
Table 37: Results obtained by using DiffPool pooling on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DiffPool 16.20 61.90 53.06 53.73
550 DiffPool 19.88 62.52 54.40 46.18
243 DiffPool 20.41 59.39 49.01 49.43
16 DiffPool 16.41 62.75 53.63 56.78
716 DiffPool 16.64 61.46 54.03 57.47
383 DiffPool 19.21 62.93 54.93 56.09
277 DiffPool 20.95 63.84 56.25 47.35
274 DiffPool 16.04 62.92 53.34 52.78
188 DiffPool 20.58 61.21 56.02 56.85
796 DiffPool 15.61 63.44 54.53 58.08
Mean $\pm$ Std 18.19 $\pm$ 2.18 62.24 $\pm$ 1.30 53.92 $\pm$ 2.02 53.47 $\pm$ 4.40
Table 38: Results obtained by using MinCut pooling on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 MinCut 22.08 70.93 63.96 46.60
550 MinCut 21.02 71.89 64.48 47.58
243 MinCut 21.43 71.52 63.83 49.56
16 MinCut 21.38 71.57 63.92 49.60
716 MinCut 21.53 71.33 61.94 47.48
383 MinCut 21.26 71.68 65.02 50.28
277 MinCut 21.09 71.80 65.17 49.75
274 MinCut 21.33 71.64 65.19 49.55
188 MinCut 21.39 71.49 64.73 50.51
796 MinCut 21.05 71.92 65.03 49.13
Mean $\pm$ Std 21.36 $\pm$ 0.31 71.58 $\pm$ 0.29 64.33 $\pm$ 0.99 49.00 $\pm$ 1.31
@@ E.4 Coauthor CS
993 DMoN 18.28 72.47 68.53 58.97
550 DMoN 18.01 72.56 67.90 57.26
243 DMoN 18.50 72.80 70.05 59.31
16 DMoN 18.04 72.93 70.70 62.98
716 DMoN 18.23 72.24 68.54 62.21
383 DMoN 17.72 73.06 70.11 60.39
277 DMoN 20.05 72.04 66.73 54.74
274 DMoN 18.13 73.24 71.11 62.35
188 DMoN 18.49 73.22 71.11 60.66
796 DMoN 20.81 71.42 67.82 53.72
Mean $\pm$ Std 18.63 $\pm$ 0.99 72.60 $\pm$ 0.58 69.26 $\pm$ 1.55 59.26 $\pm$ 3.17
@@
993 DMoN–DPR (D) 19.52 71.62 70.00 60.13
550 DMoN–DPR (D) 18.63 72.19 70.76 61.55
243 DMoN–DPR (D) 19.37 71.83 69.94 60.27
16 DMoN–DPR (D) 19.14 72.53 72.67 64.33
716 DMoN–DPR (D) 18.83 72.46 70.96 62.26
383 DMoN–DPR (D) 19.05 72.16 71.26 62.23
277 DMoN–DPR (D) 18.70 72.93 72.60 63.38
274 DMoN–DPR (D) 18.46 72.51 71.16 61.65
188 DMoN–DPR (D) 19.29 71.97 71.58 61.29
796 DMoN–DPR (D) 17.94 72.77 70.78 61.08
Mean $\pm$ Std 18.89 $\pm$ 0.48 72.30 $\pm$ 0.42 71.17 $\pm$ 0.93 61.82 $\pm$ 1.30
@@
993 DMoN–DPR (V) 18.09 72.79 69.54 60.69
550 DMoN–DPR (V) 19.20 72.05 69.23 60.66
243 DMoN–DPR (V) 17.60 73.27 71.19 62.29
16 DMoN–DPR (V) 18.45 72.43 68.76 60.54
716 DMoN–DPR (V) 18.58 72.35 68.53 59.02
383 DMoN–DPR (V) 19.05 72.25 69.61 59.12
277 DMoN–DPR (V) 19.33 72.27 69.12 57.74
274 DMoN–DPR (V) 18.62 72.97 70.38 62.66
188 DMoN–DPR (V) 18.55 73.04 71.22 60.91
796 DMoN–DPR (V) 20.01 71.56 67.98 55.77
Mean $\pm$ Std 18.75 $\pm$ 0.68 72.50 $\pm$ 0.52 69.56 $\pm$ 1.08 59.94 $\pm$ 2.08
@@
993 DMoN–DPR (E) 20.73 69.18 70.29 57.08
550 DMoN–DPR (E) 19.82 71.30 70.48 57.63
243 DMoN–DPR (E) 19.35 71.39 72.46 64.00
16 DMoN–DPR (E) 18.71 71.88 73.81 66.15
716 DMoN–DPR (E) 18.77 72.20 74.36 66.49
383 DMoN–DPR (E) 19.01 71.17 70.30 60.87
277 DMoN–DPR (E) 21.91 69.36 70.12 59.57
274 DMoN–DPR (E) 19.90 71.03 72.03 60.58
188 DMoN–DPR (E) 19.49 71.84 73.02 62.90
796 DMoN–DPR (E) 20.85 69.82 68.94 58.00
Mean $\pm$ Std 19.85 $\pm$ 1.03 70.92 $\pm$ 1.08 71.58 $\pm$ 1.81 61.33 $\pm$ 3.43
@@
993 DMoN–DPR (DV) 19.55 71.66 69.75 60.56
550 DMoN–DPR (DV) 18.24 72.37 71.14 61.70
243 DMoN–DPR (DV) 18.94 71.86 70.04 60.62
16 DMoN–DPR (DV) 19.75 71.81 72.52 63.87
716 DMoN–DPR (DV) 20.11 71.14 68.74 58.85
383 DMoN–DPR (DV) 19.52 71.72 69.75 60.47
277 DMoN–DPR (DV) 18.63 72.57 72.00 62.98
274 DMoN–DPR (DV) 18.68 72.21 71.43 62.77
188 DMoN–DPR (DV) 19.89 71.46 69.60 58.73
796 DMoN–DPR (DV) 18.13 72.81 72.20 62.90
Mean $\pm$ Std 19.14 $\pm$ 0.71 71.96 $\pm$ 0.52 70.72 $\pm$ 1.30 61.35 $\pm$ 1.78
@@
993 DMoN–DPR (DE) 19.75 70.62 71.29 63.49
550 DMoN–DPR (DE) 20.49 70.46 71.58 63.84
243 DMoN–DPR (DE) 20.83 69.25 69.19 60.23
16 DMoN–DPR (DE) 19.57 70.68 70.86 62.53
716 DMoN–DPR (DE) 19.12 71.23 71.83 62.71
383 DMoN–DPR (DE) 19.22 71.16 72.81 64.86
277 DMoN–DPR (DE) 19.78 71.78 73.49 65.51
274 DMoN–DPR (DE) 19.52 70.15 70.28 61.17
188 DMoN–DPR (DE) 20.82 69.37 67.33 57.40
796 DMoN–DPR (DE) 19.00 70.96 70.94 61.83
Mean $\pm$ Std 19.81 $\pm$ 0.68 70.57 $\pm$ 0.80 70.96 $\pm$ 1.76 62.36 $\pm$ 2.37
@@
993 DMoN–DPR (VE) 20.74 69.08 70.35 57.99
550 DMoN–DPR (VE) 19.85 71.25 70.50 57.59
243 DMoN–DPR (VE) 19.43 71.30 72.55 64.08
16 DMoN–DPR (VE) 18.74 71.85 73.81 66.17
716 DMoN–DPR (VE) 19.22 71.37 73.19 65.61
383 DMoN–DPR (VE) 18.65 71.43 70.49 60.75
277 DMoN–DPR (VE) 21.92 69.32 70.10 59.50
274 DMoN–DPR (VE) 19.93 70.96 71.96 60.58
188 DMoN–DPR (VE) 19.44 71.87 73.01 63.01
796 DMoN–DPR (VE) 20.99 69.70 68.72 57.97
Mean $\pm$ Std 19.89 $\pm$ 1.04 70.81 $\pm$ 1.04 71.47 $\pm$ 1.66 61.33 $\pm$ 3.21
@@
993 DMoN–DPR (DVE) 20.34 70.31 70.86 63.39
550 DMoN–DPR (DVE) 20.09 70.77 71.87 64.16
243 DMoN–DPR (DVE) 19.56 70.27 70.03 60.73
16 DMoN–DPR (DVE) 20.21 70.65 72.33 63.54
716 DMoN–DPR (DVE) 19.26 71.08 71.81 62.97
383 DMoN–DPR (DVE) 19.17 71.20 72.88 64.81
277 DMoN–DPR (DVE) 18.96 72.10 73.54 65.43
274 DMoN–DPR (DVE) 19.65 70.36 70.79 61.93
188 DMoN–DPR (DVE) 20.98 69.20 67.10 57.05
796 DMoN–DPR (DVE) 18.70 71.14 71.62 62.70
Mean $\pm$ Std 19.69 $\pm$ 0.71 70.71 $\pm$ 0.76 71.28 $\pm$ 1.79 62.67 $\pm$ 2.40
E.5 Coauthor Physics
The following tables list the performance of DiffPool, MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the Coauthor Physics dataset.
Table 47: Results obtained by using DiffPool pooling on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DiffPool 13.06 56.73 57.94 53.93
550 DiffPool 15.25 59.42 59.33 53.71
243 DiffPool 13.06 57.18 56.61 52.20
16 DiffPool 13.26 56.55 57.69 53.75
716 DiffPool 14.42 58.97 61.09 53.65
383 DiffPool 14.03 57.15 58.03 54.06
277 DiffPool 13.37 56.48 47.24 46.20
274 DiffPool 13.26 55.70 58.06 56.79
188 DiffPool 12.23 57.04 58.30 54.31
796 DiffPool 17.18 59.19 47.89 38.47
Mean $\pm$ Std 13.91 $\pm$ 1.42 57.44 $\pm$ 1.29 56.22 $\pm$ 4.71 51.71 $\pm$ 5.38
Table 48: Results obtained by using MinCut pooling on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 MinCut 14.08 61.78 48.91 45.03
550 MinCut 14.35 61.54 56.89 51.00
243 MinCut 13.74 61.79 51.41 47.76
16 MinCut 13.38 62.16 57.19 51.39
716 MinCut 13.18 62.55 57.10 51.27
383 MinCut 13.95 62.12 48.62 42.59
277 MinCut 14.17 61.39 48.28 46.92
274 MinCut 14.20 61.31 45.80 46.44
188 MinCut 13.33 62.51 52.14 47.82
796 MinCut 14.95 60.38 47.53 47.70
Mean $\pm$ Std 13.93 $\pm$ 0.54 61.75 $\pm$ 0.65 51.39 $\pm$ 4.30 47.79 $\pm$ 2.84